OpenStudy (anonymous):

use the quadratic formula to solve x^2+x-1=0

OpenStudy (watchmath):

Ok, try to plug in here: \[\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] where a=1,b=1 and c=-1

OpenStudy (anonymous):

hey i just did one for you, now try yourself !!!!!

OpenStudy (anonymous):

very famous number you get too!

OpenStudy (anonymous):

The equation is given in standard form, \[ax ^{2}+bx+c = 0\] the quadratic equation is what watchmath gave you. Now all you have to do is plug in a, b, and c to find what x is equal to

OpenStudy (asadkarim7):

the quadratic formula is -b +/-sqrt ( b^2-4ac)/2a substitute the value of of - b which would be -( 1) substitute the value of c which is -1 substitute the value of a which is 1 -(1) +- sqrt( (1)^- 4(1)(-1) )\2(1)

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

a = 1 b = 1 c = -1 plug in to \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

i got -1+- square root of -3

OpenStudy (anonymous):

made a mistake with the - sign in \[b^2-4ac\]

OpenStudy (anonymous):

it is \[(-1)^2-4\times 1\times -1\]

OpenStudy (anonymous):

\[=1+4=5\]

OpenStudy (anonymous):

so answer is \[\frac{-1\pm\sqrt{5}}{2}\]

OpenStudy (anonymous):

ok but i have to have two answers in fractions it wont accept it like that